OpenStudy (anonymous):

Help me please!!! Will fan and medal! What is the solution of 2x-4=16

Awolflover1 (awolflover1):

Solution for 2x-4=16 equation: 2x - 4 = 16 2x = 16 + 4 2x = 20 (divide both sides by 2 to get x) 2x/2 = 20/2 x = ?

Awolflover1 (awolflover1):

What is \[20\div2\]

OpenStudy (anonymous):

10

Awolflover1 (awolflover1):

Perfect! That is your answer. :)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

10

OpenStudy (anonymous):

I am now closing the question.